OpenStudy (anonymous):

Solve the integral (sin^2(x)*cos^2(x)) dx / (sin^8(x) - cos^8(x)) .

OpenStudy (anonymous):

is your dx supposed to be part of the numerator or you typed that wrong?

OpenStudy (anonymous):

it's regular dx, after everything

OpenStudy (anonymous):

so it's supposed to be \[(\sin ^{2}x \cos ^{2}x/\sin ^{8}x-\cos ^{8}x) dx?\]

OpenStudy (anonymous):

6 inches in diameter, 5inches in height, what is the total number of cubic inches. please help what am I doin wrong. 3.14*3squared*5* I don,t know what o do next

OpenStudy (anonymous):

presidential you can't just hijack his thread. start your own please.

OpenStudy (anonymous):

sin^8x-cos^8x is together in the denominator

OpenStudy (anonymous):

This one is kinda complicated so I'd like you to go here http://www.wolframalpha.com/input/?i=integrate+%28sin^2xcos^2x%2F%28sin^8x%E2%88%92cos^8x%29%29dx Hope that helps you out.